[Brahms-l] Shift report 20040326 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Fri Mar 26 2004 - 00:01:32 EST
------------------------------------------------------------
Date: Fri Mar 26 00:01:32 2004, Shift: 08:00-16:00
Supervisor: PS, with: 
Log book#15 pages 98 - 98

Shift Summary:
------------------------------------------------------------
 1. Two hours of beam during this shift. MC staff tried
    to refill but problem with abort kicker system occurred.
    Hasn't been fixed up to now.

 2. run 10303 - 8 deg.
    run 10305 - 8 deg., minimum bias.

 3. No detector problems.  
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Mar 26 00:01:49 2004

This archive was generated by hypermail 2.1.8 : Fri Mar 26 2004 - 00:02:13 EST